# Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation

### v6

@@ IV-D Ablation Study
As illustrated in Tab. VI, we conducted experiments on our Back2Color framework by replacing different model structures to demonstrate the effectiveness of our method. While the Depth Net learned to predict dense depth from synthetic data, it also disrupted the model’s latent memory in unsupervised learning and learned to generate sharp boundaries of obstacles. Based on the metrics in Tab. VI, there is an improvement across all metrics, particularly the $a1$ and $rmse\_log$ metrics representing detail accuracy. At the same time, we also observed that there was no significant difference between only using SHIFT and vKitti2. In Tab. VI, we also tested training VADepth from scratch without ImageNet pre-training. The results demonstrated that using a pre-trained checkpoint enhances the network’s ability to process natural images, significantly benefiting depth estimation.
@@
TABLE VI: Experiments on the Kitti Eigen Split at 192×640 resolution with different models for Back2Color.VAN* means not apply ImageNet pretrained checkpoints.
@@
Additionally, we used various synthetic datasets such as vKitti2, SHIFT, and UrbanSyn to real-world unsupervised depth model training, as shown in Tab.VII. The results consistently achieved excellent improvement, demonstrating the strong adaptability of our method. To demonstrate that the improvement in our Back2Color method is not due to the increased training samples (Synthetic Datasets) but rather due to the diverse priors brought by various synthetic datasets, we conducted experiments using 20% and 50% synthetic samples (Rows 6, 7, and Rows 9, 10) while ensuring the total number of dataset samples remains the same as the KITTI dataset (39,810). The results are consistent with those obtained using the full dataset, proving that the enhancement of our method is not due to the increase in training samples. Additionally, when using the MixSyn mixed synthetic dataset, some metrics (rmse_log and a1) show significant improvement compared to using a single dataset, demonstrating that our method benefits from learning the variety of multiple synthetic datasets.
@@
TABLE VII: Ablation Study on the impact of Various Synthetic Datasets.
@@
192x640 vKitti2 39810 21260 0 0 0.094 0.694 4.296 0.173 0.907 0.968 0.984
192x640 SHIFT 39810 0 153000 0 0.092 0.676 4.238 0.171 0.909 0.969 0.984
192x640 Urban 39810 0 0 7539 0.093 0.644 4.198 0.172 0.907 0.968 0.984
192x640 MixSyn 39810 21260 153000 7539 0.092 0.685 4.229 0.170 0.911 0.968 0.984
320x1024 vKitti2 39810 21260 0 0 0.090 0.677 4.173 0.169 0.914 0.969 0.984
320x1024 vKitti2 31848 7962 0 0 0.090 0.675 4.143 0.168 0.917 0.970 0.985
320x1024 vKitti2 19905 19905 0 0 0.091 0.684 4.192 0.169 0.913 0.969 0.985
320x1024 SHIFT 39810 0 153000 0 0.089 0.651 4.158 0.167 0.916 0.970 0.985
320x1024 SHIFT 31848 0 7962 0 0.090 0.684 4.158 0.168 0.916 0.970 0.984
320x1024 SHIFT 19905 0 19905 0 0.090 0.680 4.144 0.167 0.915 0.970 0.985
320x1024 Urban 39810 0 0 7539 0.090 0.654 4.084 0.167 0.915 0.970 0.985
320x1024 MixSyn 39810 21260 153000 7539 0.087 0.644 4.079 0.166 0.920 0.970 0.985
@@
TABLE VIII: Ablation study on models with params and computational complexities(GFLOPs).
@@
MonoViT[2] small 320x1024 33.63 159.14 0.093 0.671 4.202 0.169 0.912 0.969 0.985
Lite-mono[1] 8M 320x1024 8.76 29.88 0.097 0.710 4.309 0.174 0.905 0.967 0.984
Back2Color VAN0 320x1024 5.17 17.64 0.094 0.706 4.257 0.171 0.910 0.969 0.984
Back2Color VAN1 320x1024 18.61 51.48 0.090 0.664 4.139 0.168 0.915 0.970 0.984
Back2Color VAN2 320x1024 31.32 76.01 0.087 0.644 4.079 0.166 0.920 0.970 0.985
Back2Color VAN3 320x1024 49.5 114.63 0.087 0.653 4.068 0.165 0.919 0.971 0.985
Back2Color VAN0 384x1280 5.17 17.64 0.093 0.676 4.169 0.171 0.911 0.969 0.984
Back2Color VAN1 384x1280 18.61 51.48 0.090 0.664 4.102 0.168 0.917 0.970 0.984
Back2Color VAN2 384x1280 31.32 76.01 0.086 0.629 4.054 0.165 0.920 0.970 0.985
Back2Color VAN3 384x1280 49.5 114.63 0.087 0.652 4.062 0.165 0.920 0.970 0.984
@@ IV-F Discussion
IV-F1 Why can the network transform information beyond depth
@@
IV-F2 Why learn Depth and Color Net simultaneously online
the primary objective of our method is depth prediction. As the Depth Net undergoes continuous updates to enhance its accuracy and performance, it is imperative to ensure that the Color Net also adapts accordingly. This synchronization is crucial for achieving superior color transformation effects. By reprojecting the Color Net’s adjustments with the updates in the Depth Net, we aim to optimize the overall performance of the model. This dynamic adjustment process ensures that the color transformation remains effective and reproject with the evolving depth predictions.
IV-F3 Can the VAN-based encoder replace the ResNet structure in the Pose encoder
so far, we have not observed any new research that can replace the ResNet-based PoseEncoder[9], whether using CNN-based or Transformer-based methods. We have attempted some alternatives, but none have improved depth estimation performance; in fact, many have resulted in performance degradation. In this study, we also tried using a VAN-based model to replace the ResNet-based PoseEncoder, but still did not observe any significant improvement. We believe that the pose estimation task is a small parameter regression task that is relatively simple compared to depth prediction and does not require a powerful network.
@@ V Conclusion
In this paper, we presented Back to the Color(Back2Color), a novel framework designed to enhance unsupervised depth estimation by leveraging synthetic data. Our method addresses the challenge of aligning synthetic world with real world colors, enabling the transformation of predicted synthetic depth into realistic colors with high fidelity. By joint training on synthetic and real-world datasets using Back2Color, we observed substantial improvements in depth estimation performance. Additionally, we designed a depth net based on the Visual Attention Network (VAN), which allowed our CNN-based network to surpass several Transformer-based depth models. Furthermore, we improved the uncertainty-based loss fusion method by combining the advantages of temporal and spatial reprojection, thoroughly addressing issues related to non-rigid motion. Experimental results demonstrated the effectiveness of our method on both the KITTI and Cityscapes datasets. Overall, Back2Color offers a promising solution for advancing unsupervised depth estimation by effectively integrating synthetic data into the training process.

### v7

@@ IV-D Ablation Study
TABLE VI: Ablation experiments to test different modules. The experiments use a stereo setup(ms) with a resolution of 320×1024. Row 1 shows the results of Monodepth2[14] as a reference.
Modules Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
Backbone 0.094 0.703 4.435 0.173 0.909 0.969 0.984
Backbone+MixSyn 0.095 0.724 4.672 0.172 0.910 0.969 0.984
Backbone+MixSyn+B2C 0.090 0.688 4.320 0.169 0.915 0.970 0.985
Backbone+MixSyn+B2C+UTSF 0.087 0.644 4.079 0.166 0.920 0.970 0.985
As illustrated in Tab. VII, we conducted experiments on our Back2Color framework by replacing different network structures to demonstrate the effectiveness of our new framework. The results also demonstrated that using a pre-trained checkpoint enhances the network’s ability to process natural images, significantly benefiting depth estimation. Next, we conducted ablation experiments to separately evaluate each of the proposed modules, ad shown in Tab.VI including the MixSyn dataset, the Back2Color training framework, and the auto-learning uncertainty temporal-spatial fusion method (UTSF). The results demonstrated that simply combining real and synthetic data for training did not improve performance. However, applying the Back2Color framework for color domain adaptation effectively addressed this issue. Additionally, Auto-UTSF further improved performance by leveraging advantageous information from both the temporal and spatial domains.
@@
TABLE VII: Experiments on the KITTI Eigen split at 192×640 resolution. Results show that Back2Color consistently improves performance across different backbones, indicating that the gain is not solely attributed to backbone selection. VAN* denotes training without ImageNet pretraining.
@@
Additionally, we used various synthetic datasets such as vKITTI2, SHIFT, and UrbanSyn to real-world unsupervised depth model training, as shown in Tab.VIII. The results consistently achieved excellent improvement, demonstrating the strong adaptability of our method. To demonstrate that the improvement in our Back2Color method is not due to the increased training samples (Synthetic Datasets) but rather due to the diverse priors brought by various synthetic datasets, we conducted experiments using 20% and 50% synthetic samples (Rows 6, 7, and Rows 9, 10) while ensuring the total number of dataset samples remains the same as the KITTI dataset (39,810). The results are consistent with those obtained using the full dataset, proving that the enhancement of our method is not due to the increase in training samples. Additionally, when using the MixSyn mixed synthetic dataset, some metrics (rmse_log and a1) show significant improvement compared to using a single dataset, demonstrating that our method benefits from learning the variety of multiple synthetic datasets.
@@
TABLE VIII: Ablation Study on the impact of Various Synthetic Datasets.
@@
[…]
@@
TABLE IX: Ablation study on models with params and computational complexities(GFLOPs).
@@
[…]
@@ IV-F Discussion
Why can the network transform information beyond depth?
@@
In practice, the ambiguity is significantly reduced by the strong appearance consistency within a target domain (e.g., KITTI or Cityscapes), enabling the depth-to-color network to learn a stable, domain-specific mapping.
@@ V Conclusion
This paper presented Back2Color, an unsupervised monocular depth estimation framework for robust depth perception in urban driving scenarios. By learning a depth-to-color mapping from real-world data and transferring it to synthetic scenes, the proposed method effectively reduces the synthetic-to-real appearance gap and improves depth estimation under real traffic conditions. In addition, an uncertainty-aware temporal-spatial fusion strategy is introduced to handle non-rigid motion and occlusions, leading to more reliable depth predictions in dynamic environments. Experimental results on standard driving benchmarks demonstrate consistent performance improvements with favorable computational efficiency, highlighting the practicality of the proposed approach for vision-based perception in intelligent transportation systems.